OpenStudy (anonymous):

I think it's A? Am I right? Which equation describes the line that contains (1, 5) and has a slope of 2? a. y = 2x + 3 b. y = 2x - 3 c. y = 2x - 9 d. y = 2x + 9

jimthompson5910 (jim_thompson5910):

you are correct

jimthompson5910 (jim_thompson5910):

check: y = 2x + 3 ... this has a slope of 2 5 = 2(1) + 3 ... plug in (1,5) 5 = 2 + 3 5 = 5

jimthompson5910 (jim_thompson5910):

so that confirms the answer

OpenStudy (anonymous):

Thanks.

jimthompson5910 (jim_thompson5910):

np